OpenStudy (anonymous):

What is the simplified expression for 3 to the power of negative 4 multiplied by 2 to the power of 3 multiplied by 3 to the power of 2 whole over 2 to the power of 4 multiplied by 3 to the power of negative 3? 3 over 2 3 to the power of 2 over 2 to the power of 2 3 to the power of 2 over 2 2 to the power of 4 over 3

OpenStudy (anonymous):

@mathmale @surjithayer

OpenStudy (anonymous):

@THEHELPER123

OpenStudy (anonymous):

\[\frac{ x^a*y^b*z^c }{ y^f*x^g*z^h }=x ^{(a-g)}*y ^{\left( b-f \right)}*z ^{\left( c-h \right)}\]